# FaceTalk: Audio-Driven Motion Diffusion for Neural Parametric Head Models

### 13 Dec 2023 v1

@@ 8 Acknowledgments
We provide additional ablation studies regarding dataset creation and model architecture in Section A, network architecture details in Section  B, metric evaluation in Section C, and additional details regarding dataset creation in Section  D. For a visual comparisons of all experiments and ablation studies, we request readers to watch the supplemental video.
@@ A Additional Ablation Studies
Our Method LSE-D $\downarrow$ Diversity $\uparrow$ FID $\downarrow$ KID $\downarrow$
Input 12.029 0.006 47.168 0.0095
Scale-Shift 11.9575 0.204 42.769 0.0093
SqueezeExcite [33] 11.7586 0.029 41.028 0.0092
Dot Prod. Attention 12.83 2.009e-7 121.584 0.08
FiLM [50] 11.2737 0.340 40.692 0.0092
@@ B Architecture and Training Details
Our expression decoder consists of a stack of transformer-based decoder blocks with FiLM [50] layers sandwiched between Multihead Self-Attention, Multihead Cross-Attention and Feedforward Layers, as shown in Fig. 10(c). We use a stacked multi-head transformer decoder model with latent dimension of 256. For Multihead Self- and Cross-Attention layers, we use eight heads and set the dimension to 1024 for each transformer decoder block. The linear projection layer at input projects from the NPHM expression space (200-dim.) to the latent space of our expression decoder (256-dim.), and processes them through the stack of decoder blocks and finally use another linear layer to project the learnt expressions from our latent space (256-dim.) back to NPHM expression space (200-dim).

### 17 Mar 2024 v2

@@ 8 Acknowledgments
We provide additional ablation studies regarding flame baselines, landmark conditioning, dataset creation and model architecture in Section A, network architecture details in Section  B, metric evaluation in Section C, and additional details regarding dataset creation in Section  D. For a visual comparisons of all experiments and ablation studies, we request readers to watch the supplemental video.
@@ A Additional Ablation Studies
Baselines trained on FLAME fittings of our data.
We report results on our hybrid test set with audio from Vocaset, Ours, and LJSpeech. We additionally train baselines on our dataset (using FLAME fittings, as they don’t operate on NPHM space) and show results on the same hybrid test set in Tab. 3. Emote and EmoTalk require emotional correspondences and MeshTalk is incompatible with the FLAME topology; thus, a comparison with them is not possible. Our method outperforms baselines on lip-sync (lowest LSE-D), while significantly improving quality (highest FIQA/VQA).
Method LSE-D $\downarrow$ FIQA $\uparrow$ VQA $\uparrow$
FaceFormer 12.1391 38.64 0.4206
CodeTalker 11.8442 39.44 0.4218
Imitator 11.7402 37.83 0.4492
FaceDiffuser 12.7644 37.65 0.3826
Ours (Flame) 11.6839 38.53 0.4868
Ours (w/o diffusion) 11.3217 42.81 0.5469
Ours (Full) 11.2737 45.75 0.6145
Table 3: All methods, including baselines were trained on our dataset. The baselines and Ours (Flame) were trained with FLAME fittings, and bottom two rows with NPHM fittings. Our Flame baseline also outperforms other methods in lip sync and video quality. Our full model synthesizes highest-quality animations (highest VQA), well-matches real face qualities (highest FIQA), while performing better lip sync (lowest LSE-D).
Additional Conditioning.
@@
Our Method LSE-D $\downarrow$ Diversity $\uparrow$
Input 12.029 0.006
Scale-Shift 11.9575 0.204
SqueezeExcite [34] 11.7586 0.029
Dot Prod. Attention 12.83 2.009e-7
FiLM [52] 11.2737 0.340
@@ B Architecture and Training Details
Our expression decoder consists of a stack of four transformer-based decoder blocks with FiLM [52] layers sandwiched between Multihead Self-Attention, Multihead Cross-Attention and Feedforward Layers, as shown in Fig. 10(c). We use a stacked multi-head transformer decoder model with latent dimension of 256. For Multihead Self- and Cross-Attention layers, we use eight heads and set the dimension to 1024 for each transformer decoder block. The linear projection layer at input projects from the NPHM expression space (200-dim.) to the latent space of our expression decoder (256-dim.), and processes them through the stack of decoder blocks and finally use another linear layer to project the learnt expressions from our latent space (256-dim.) back to NPHM expression space (200-dim).